Write the equation for the problem and solve. The product of two consecutive even integers is equal to 120. What are all the integers? No credit for guesses or interpolation.
Answer by mananth(16946)

let one number be x
the other number will be x+2
..
x(x+2)= 120
x^2+2x=120
x^2+2x-120=0
x^2+12x-10x-120=0
x(x+12)-10(x+12)
(x+12)(x-10)
x=-12 OR x=10
If x=-12
then the numbers are -12, -10
If x=10 Then the numbers are 10, 12